OpenStudy (janewelcomes):

Find all the zeros of the equation. -3x^4+27x^2+1200=0

OpenStudy (cwrw238):

first simplify by dividing through by 3

OpenStudy (cwrw238):

in fact divide by -3

OpenStudy (janewelcomes):

So would it be -x^4+9x^2+400=0?

OpenStudy (cwrw238):

which makes it easier to factor

OpenStudy (cwrw238):

x^4 - 9x^2 - 400 = 0

OpenStudy (janewelcomes):

Oooh. Ok. I divided by 3 instead of -3.

OpenStudy (cwrw238):

now this will factor

OpenStudy (janewelcomes):

How do I factor it?

OpenStudy (cwrw238):

it will be of the form (x^2 + __)(x^2 - __)

OpenStudy (cwrw238):

you need to find 2 numbers whose product is -400 and whose sum = -9

OpenStudy (janewelcomes):

Two numbers who sum to -9 are -4 and -5. I don't remember how to do this. I'm sorry. I feel really stupid right now.

OpenStudy (cwrw238):

factor each will help 400 = 2*2*2*2*5*5 9 = 3*3 -400 = 16 *-25 and 16 - 25 = -9

OpenStudy (cwrw238):

so we have (x^2 - 25)(x^2+16) = 0

OpenStudy (janewelcomes):

That's just what I was getting ready to say. YAY! I'm feeling a little smarter now.

OpenStudy (cwrw238):

so now you need to solve x^2-25 = 0 and x^2 -16 = 0

OpenStudy (cwrw238):

as its a quartic equation there will be 4 zeros 2 will be real and 2 complex

OpenStudy (janewelcomes):

Right now I have X^2=25 and x^2=16. Now I divide by 2 right? That way I have just x= instead of x^2= right?

OpenStudy (cwrw238):

sorry - i had one equation wrong irs x^2 + 16 = 0

OpenStudy (cwrw238):

no you take the square root of each side

OpenStudy (janewelcomes):

Ooooh! Duh. Thank you. Sorry.

OpenStudy (cwrw238):

x^2 = 25 x = +/- sqrt25 = +5, -5

OpenStudy (janewelcomes):

Oh! Got it! So the zeros of the equation are 5,-5, 4i, and -4i right?

OpenStudy (cwrw238):

right

OpenStudy (janewelcomes):

Yay! Thank you so much! How do I give you a medal?

OpenStudy (cwrw238):

click on the blue box Best Response

OpenStudy (janewelcomes):

Done. Thanks again!